# N2: A Unified Python Package and Test Bench for Nearest Neighbor-Based Matrix Completion

### v1

@@ 1 Introduction
Nearest neighbor methods are a class of non-parametric algorithms widely used for regression, classification and pattern recognition. Due to their scalability and success under models with minimal assumptions, nearest neighbor methods have recently been adopted for practical fields such as matrix completion and counterfactual inference in panel data settings. Matrix completion is a well-established field that supplies practitioners with many tools to recover underlying matrices using partial or even noisy observations HMLZ (15); Cha (15); KMO (10), with recommendation systems KBV (09); Rec (11) as an important use-case. Panel data counterfactual inference aims at learning the treatment effect of policies across time Bai (09); BN (21); ABD+ (21). One important example is individualized healthcare predictions KSS+ (19). Nearest neighbor methods were recently recognized as effective in providing granular inference guarantees for both matrix completion and counterfactual inference when either the missingness or the policy treatment are not completely random and confounded MC (19); DTT+22a ; ADSS (23). They have also been recently leveraged to tackle distributional matrix completion settings FCAD (24); CFC+ (24)
Despite nearest neighbor methods popularity, there is no unified package that lets a user easily switch between different kinds of nearest neighbor algorithms for matrix completion and counterfactual inference. In this paper, we present a package ($\mathbf{N}^{2}$) to unify several nearest neighbor methods under a single interface, so users can easily choose the method that suits their data the best, for both scalar and distributional missing data problems. Next, we review two broad applications of nearest neighbors and matrix completion: recommendation systems and panel data settings.
@@
In the problem of a recommendation system, multiple users submit ratings / preferences for a subset of items, and the vendor recommends items based on the partial rating information RS (05). Prediction of unseen ratings is the key to a good recommendation, and the predictions are based on the patterns of the partially observed ratings which can be formalized into a matrix completion problem. A popular instance is the Netflix problem, an open competition for the best prediction algorithm of user ratings of films. Users (rows of matrix) have the option to rate the films (columns of matrix). However, users typically rate only a subset of films, and the ratings are commonly biased towards their preferences. Under the common belief that only a few factors drive the preference and ratings of the users, many matrix completion algorithms have resorted to the low rank assumption on the ratings matrix Rec (11); CR (12). Variants of nearest neighbors (in short, $\mathsf{NN}$) recently were introduced to correct for the potential bias originating from the fact that observation patterns are influenced by the matrix entries (e.g. ratings) themselves and by unmeasured confounders ADSS (23); AADS (24).
@@
Panel data constitutes measurements of multiple units that are tracked over time (hence a type of longitudinal data) as they undergo different types of interventions at different times Bai (09). Panel data is used to analyze the causal effects of any policies, treatments, or business decisions that affect the subjects in the data, making this data type essential in econometrics and health studies. Matrix completion algorithms enable estimation of these causal effects at the unit×$\times$×time level, the most granular scale where causal effects can be estimated ABD+ (21); ADSS (23). For example, consider a mobile health app trying to learn the effectiveness of two exercise routines. Suppose the app alternates two routines across $N$ different users repeatedly over $T$ weeks, where their health activities (say physical step counts) throughout each week are recorded. Then, we can use matrix completion to impute the counterfactual step counts under treatment and control, allowing us to estimate causal effects at the most granular scale.
@@
We make the following contributions: First, we present a unified framework for nearest neighbor algorithms that facilitates extending to new variants, especially for matrix completion problems. We leverage this framework to introduce then a new nearest neighbors algorithm, auto nearest neighbors (in short $\mathsf{AutoNN}$), which aims both to improve on the existing methods and demonstrate the ease of developing new variants with our library. Next, we introduce a unified, easy-to-implement nearest neighbor library that contains a breadth of nearest neighbor algorithms for matrix comple2tion problems. Next, we present a test bench called $\mathbf{N}^{2}$-Bench that comprises of several real-world data sets across applications. Finally, we illustrate our library’s wide applicability on the test benchmark.
@@ 1.1 Related work
As introduced above, nearest neighbor methods are widely used for non-parametric regression, classification, and pattern recognition CH (67); CBS (20). Recently, nearest neighbor methods were introduced as an effective algorithm for matrix completion problems LSSY (19); ADSS (23); DTT+22a , especially when the missingness depended on observations and unobserved confounders. The fact that nearest neighbor methods target a single entry at a time via matching makes them effective against various types of missing patterns. The class of algorithms has grown to account for a wide range of applications involving scalar or distributional settings; for instance, nearest neighbors are used for evaluating the personalized recommendation effect of a healthcare app on the average and the distribution of physical step counts DTT+22a ; CFC+ (24); FCAD (24); SPD (24) and are also used for policy evaluations ADSS (23). Our library, $\mathbf{N}^{2}$, is designed to easily implement the vanilla scalar versions of nearest neighbors LSSY (19); DTT+22a as well as its unweighted DTT+22b ; SPD (24) and weighted SPD (25) variants. Finally, $\mathbf{N}^{2}$ is also capable of distributional matrix completion CFC+ (24); FCAD (24).
@@
Universal singular value thresholding (USVT), proposed in Cha (15); BC (22), is a classical spectral-based method for performing matrix completion; its core functionality is based on a singular value decomposition of the matrix and thresholding the singular values. SoftImpute, introduced by HMLZ (15), is another widely used optimization-based algorithm for matrix completion. The algorithm computes ridge-regression updates of the low-rank factors iteratively and finally soft-thresholds the singular values to impose a nuclear norm penalty. Notably USVT and SoftImpute have provable guarantees when missingness is completely at random, but empirically fail when the missing pattern depends on the observed entries or the unobserved confounders ADSS (23). Our real-world analysis in Sec. 4 once again demonstrates this point.
@@
Scikit-Learn PVG+ (11), a popular Python package for machine learning tools, implements a simple k-nearest neighbor algorithm for imputing missing values in a feature matrix. However, their implementation is designed for the feature matrix setting. So, neighbors are only defined across samples (row-wise). Additionally, they do not provide any implementation for more advanced nearest neighbor algorithms, nor does their package allow for easy extendability like our proposed package.
Organization
Sec. 2 contains an overview of the existing nearest neighbor algorithms implemented in our library, $\mathbf{N}^{2}$. Sec. 2.3 introduces a new nearest neighbor algorithm, $\mathsf{AutoNN}$, that is also implemented in $\mathbf{N}^{2}$. Sec. 3 touches upon the high-level (class) structure of $\mathbf{N}^{2}$, as well as the interface for practitioners. Sec. 4 tests different variants of nearest neighbor methods and classical methods on our new test bench of diverse datasets called $\mathbf{N}^{2}$-Bench. Finally, in Sec. 5, we provide concluding remarks and outline future directions of research.
@@ 2 Nearest Neighbors for Matrix Completion
[…]
@@ 2.1 Unified framework
Let $\varphi(x,x^{\prime})$ be a metric between $x,x^{\prime}\in\mathcal{X}$ for some space $\mathcal{X}$. Further define $\widehat{\varphi}(Z_{i,t},Z_{j,s})$ as a data-dependent distance between any two observed entries $(i,t)$ and $(j,s)$ of the matrix 1. The two modules can now be defined:
@@
    $\widehat{\theta}=\mathop{\mathrm{missing}}{argmin}_{x\in\mathcal{X}}\sum_{j\in[N],s\in[T]}w_{j,s}A_{j,s}\varphi(x,Z_{j,s}).$ (4)
The Distance module calculates the row-wise and column-wise distance of the matrix, by taking the average of the observed entry-wise distance $\widehat{\varphi}(\cdot,\cdot)$. The Average module calculates the weighted average of observed measurements, where the notion of average depends on the metric $\varphi$ and the space $\mathcal{X}$ on which the metric $\varphi$ is defined. Notably, the weights $\mathcal{W}$ in the Average module encodes the entry information of the estimand.
@@
[…]
@@ 2.2 Existing methods
We present existing variants of nearest neighbors using the two modules introduced Sec. 2.1; all the methods presented here are recovered by sequentially applying Distance and Average with the appropriate specification of $\widehat{\varphi}$, $\varphi$ and $\mathcal{W}$.
All methods except $\mathsf{AWNN}$ and our newly proposed $\mathsf{AutoNN}$, have binary weights i.e., $w_{j,s}\in\{0,1\}$. $\mathsf{AutoNN}$, detailed in Sec. 2.3, uses weights to carefully pool together the benefits of $\mathsf{TSNN}$ and $\mathsf{DRNN}$. $\mathsf{AWNN}$ SPD (25) improves upon $\mathsf{RowNN}$ by adaptively choosing the weights which optimally balances the bias-variance tradeoff of $\mathsf{RowNN}$ as follows
$\big{(}w_{1}^{\star}(i,t),...,w_{N}^{\star}(i,t)\big{)}:=\mathop{\mathrm{missing}}{argmin}_{(v_{1},...,v_{N})\in\Delta_{N}}2\log(2N)\widehat{\sigma}^{2}\sum_{k\in[N]}v_{k}^{2}+\sum_{k\in[N]}v_{k}A_{k,t}\rho_{i,k}^{\mathrm{row}}.$ (5)
where $\widehat{\sigma}^{2}$ is the estimated error and $\Delta_{N}$ is a simplex in N; see SPD (25) for details of 5. Tab. 1 contains a concise summary of the existing nearest neighbor variants; see App. B for a detailed exposition for each methods.
@@
$n=1$ $\mathsf{RowNN}$  (LSSY, 19) (Alg. 1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{ColNN}$  (LSSY, 19) (Alg. 1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq 0,\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{TSNN}$  (SPD, 24) (Alg. 2) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{AWNN}$  (SPD, 25) (Alg. 5) $(x-y)^{2}$ $(x-y)^{2}$ $w_{j}^{\star}(i,t)\cdot\mathbf{1}({\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{DRNN}$  (DTT+22b, ) (Alg. 3) $\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$
$\mathsf{AutoNN}$  (Sec. 2.3) $\alpha\cdot\mathsf{DRNN}+(1-\alpha)\cdot\mathsf{TSNN}$
$n>1$ $\mathsf{KernelNN}$  (CFC+, 24) (Alg. 4) $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(x,y)$ $\mathsf{MMD}_{\mathbf{k}}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{W_{2}NN}$  (FCAD, 24) (Alg. 4) $\widehat{\mathsf{W}}_{2}^{2}(x,y)$ $\mathsf{W}_{2}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
Under the distributional matrix completion setting ($n>1$ in 1), the methods $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ in Tab. 1 take $\mu,\nu\in\mathcal{X}$ as square integrable probability measures, and $\varphi(\mu,\nu)$ as either the squared maximum mean discrepency (i.e. $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu,\nu)$, see MFS+ (17)) or squared Wasserstein metric (i.e., $\mathsf{W}_{2}(\mu,\nu)$, see Big (20)). Further, the entry-wise distance $\widehat{\varphi}(x,y)$ in this case is either the unbiased U-statistics estimator $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu_{i,t},\mu_{j,s})$ (see MFS+ (17)) or the quantile based estimator $\widehat{\mathsf{W}}_{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{W}_{2}(\mu_{i,t},\mu_{j,s})$ (see Big (20)).

### v2

@@ 1 Introduction
Nearest neighbor methods are a class of non-parametric algorithms widely used for regression, classification and pattern recognition. Due to their scalability and success under models with minimal assumptions, nearest neighbor methods have recently been adopted for practical fields such as matrix completion and counterfactual inference in panel data settings. Matrix completion is a well-established field that supplies practitioners with many tools to recover underlying matrices using partial or even noisy observations hastie2015matrix; chatterjee2015matrix; keshavan2010matrix, with recommendation systems koren2009matrix; recht2011simpler as an important use-case. Panel data counterfactual inference aims at learning the treatment effect of policies across time bai2009panel; bai2021matrix; athey2021matrix. One important example is individualized healthcare predictions klasnja2019efficacy. Nearest neighbor methods were recently recognized as effective in providing granular inference guarantees for both matrix completion and counterfactual inference when either the missingness or the policy treatment are not completely random and confounded ma2019missing; dwivedi2022counterfactual; agarwal2023causal. They have also been recently leveraged to tackle distributional matrix completion settings feitelberg2024distributional; choi2024learning
Despite nearest neighbor methods popularity, there is no unified package that lets a user easily switch between different kinds of nearest neighbor algorithms for matrix completion and counterfactual inference. In this paper, we present a package [^1] https://github.com/aashish-khub/NearestNeighbors to unify several nearest neighbor methods under a single interface, so users can easily choose the method that suits their data the best, for both scalar and distributional missing data problems. Next, we review two broad applications of nearest neighbors and matrix completion: recommendation systems and panel data settings.
@@
In the problem of a recommendation system, multiple users submit ratings / preferences for a subset of items, and the vendor recommends items based on the partial rating information rennie2005fast. Prediction of unseen ratings is the key to a good recommendation, and the predictions are based on the patterns of the partially observed ratings which can be formalized into a matrix completion problem. A popular instance is the Netflix problem, an open competition for the best prediction algorithm of user ratings of films. Users (rows of matrix) have the option to rate the films (columns of matrix). However, users typically rate only a subset of films, and the ratings are commonly biased towards their preferences. Under the common belief that only a few factors drive the preference and ratings of the users, many matrix completion algorithms have resorted to the low rank assumption on the ratings matrix recht2011simpler; candes2012exact. Variants of nearest neighbors (in short, $\mathsf{NN}$) recently were introduced to correct for the potential bias originating from the fact that observation patterns are influenced by the matrix entries (e.g. ratings) themselves and by unmeasured confounders agarwal2023causal; abadie2024doubly.
@@
[…]
@@
We make the following contributions: First, we present a unified framework for nearest neighbor algorithms that facilitates extending to new variants, especially for matrix completion problems. We leverage this framework to introduce then a new nearest neighbors algorithm, auto nearest neighbors (in short $\mathsf{AutoNN}$), which aims both to improve on the existing methods and demonstrate the ease of developing new variants with our library. Next, we introduce a unified, easy-to-implement nearest neighbor library that contains a breadth of nearest neighbor algorithms for matrix completion problems. Next, we present a test bench called $\mathbf{N}^{2}$-Bench that comprises of several real-world data sets across applications. Finally, we illustrate our library’s wide applicability on the test benchmark.
@@ 1.1 Related work
As introduced above, nearest neighbor methods are widely used for non-parametric regression, classification, and pattern recognition cover1967nearest; cannings2020local. Recently, nearest neighbor methods were introduced as an effective algorithm for matrix completion problems li2019nearest; agarwal2023causal; dwivedi2022counterfactual, especially when the missingness depended on observations and unobserved confounders. The fact that nearest neighbor methods target a single entry at a time via matching makes them effective against various types of missing patterns. The class of algorithms has grown to account for a wide range of applications involving scalar or distributional settings; for instance, nearest neighbors are used for evaluating the personalized recommendation effect of a healthcare app on the average and the distribution of physical step counts dwivedi2022counterfactual; choi2024learning; feitelberg2024distributional; sadhukhan2024adaptivity and are also used for policy evaluations agarwal2023causal. Our library, $\mathbf{N}^{2}$, is designed to easily implement the vanilla scalar versions of nearest neighbors li2019nearest; dwivedi2022counterfactual as well as its unweighted dwivedi2022doubly; sadhukhan2024adaptivity and weighted sadhukhan2025adaptivelyweightednearestneighborsmatrix variants. Finally, $\mathbf{N}^{2}$ is also capable of distributional matrix completion choi2024learning; feitelberg2024distributional.
@@
Universal singular value thresholding (USVT), proposed in chatterjee2015matrix; bhattacharya2022matrix, is a classical spectral-based method for performing matrix completion; its core functionality is based on a singular value decomposition of the matrix and thresholding the singular values. SoftImpute, introduced by hastie2015matrix, is another widely used optimization-based algorithm for matrix completion. The algorithm computes ridge-regression updates of the low-rank factors iteratively and finally soft-thresholds the singular values to impose a nuclear norm penalty. Notably USVT and SoftImpute have provable guarantees when missingness is completely at random, but empirically fail when the missing pattern depends on the observed entries or the unobserved confounders agarwal2023causal. Our real-world analysis in Sec.˜4 once again demonstrates this point.
@@
Scikit-Learn scikit-learn, a popular Python package for machine learning tools, implements a simple k-nearest neighbor algorithm for imputing missing values in a feature matrix. However, their implementation is designed for the feature matrix setting. So, neighbors are only defined across samples (row-wise). Additionally, they do not provide any implementation for more advanced nearest neighbor algorithms, nor does their package allow for easy extendability like our proposed package.
@@ 2 Nearest Neighbors for Matrix Completion
In other words, for matrix entries where $A_{i,t}=1$, we observe $n$ measurements $Z_{i,t}$ that takes value $X_{i,t}$ realized from distribution $\mu_{i,t}$. When $n=1$, i.e., $Z_{i,t}=X_{1}(i,t)$, we refer to ˜2 as the scalar matrix completion model; scalar matrix completion is the most common problem posed in the literature candes2012exact; recht2011simpler; koren2009matrix; hastie2015matrix; chatterjee2015matrix; dwivedi2022counterfactual; dwivedi2022doubly; agarwal2023causal, where the goal is to learn the mean of the underlying distributions $\{\theta_{i,t}=\int xd\mu_{i,t}(x)\}_{i\in[N],t\in[T]}.$ When there are more than one observed measurements per entry, i.e., $Z_{i,t}=[X_{1}(i,t),...,X_{n}(i,t)]$ for $n\geq 2$, we refer to ˜2 as the distributional matrix completion problem, the goal being the recovery of the distributions as a whole. We refer the readers to App.˜A for a detailed discussion on the structural assumptions imposed on the model ˜2.
@@ 2.1 Unified framework
Let $\varphi(x,x^{\prime})$ be a metric between $x,x^{\prime}\in\mathcal{X}$ for some space $\mathcal{X}$. Further define $\widehat{\varphi}(Z_{i,t},Z_{j,s})$ as a data-dependent distance between any two observed entries $(i,t)$ and $(j,s)$ of the matrix ˜2. The two modules can now be defined:
@@
    $\widehat{\theta}=\mathop{\mathrm{argmin}}_{x\in\mathcal{X}}\sum_{j\in[N],s\in[T]}w_{j,s}A_{j,s}\varphi(x,Z_{j,s}).$ (5)
[…]
@@
The vanilla row-wise nearest neighbors li2019nearest that targets the mean $\theta_{i,t}=\int xd\mu_{i,t}(x)$ of entry $(i,t)$ is recovered by first applying Distance with $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})={(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}}$, applying Average with the non-smooth weight $w_{j,s}=\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1}})\cdot\mathbf{1}({\rho_{s,t}^{\mathrm{col}}\leq 0})$}, and using the metric $\varphi(x,y)=(x-y)^{2}$. Note that the non-smooth weight satisfies $w_{j,t}=\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1}})$, whereas $w_{j,s}=0$ for $s\neq t$; by defining the nearest neighbor set $\mathbf{N}_{t,\eta_{1}}:=\{j\in[N]:\rho_{j,t}^{\mathrm{row}}\leq\eta_{1}\}$, the Average module output can be rewritten as $\mathop{\mathrm{argmin}}_{x\in}\sum_{j\in\mathbf{N}_{t,\eta_{1}}}A_{j,t}(x-Z_{j,t})^{2}=|\mathbf{N}_{t,\eta_{1}}|^{-1}\sum_{j\in\mathbf{N}_{t,\eta_{1}}}Z_{j,t}$}.
@@ 2.2 Existing methods
We present existing variants of nearest neighbors using the two modules introduced Sec.˜2.1; all the methods presented here are recovered by sequentially applying Distance and Average with the appropriate specification of $\widehat{\varphi}$, $\varphi$ and $\mathcal{W}$.
All methods except $\mathsf{AWNN}$ and our newly proposed $\mathsf{AutoNN}$, have binary weights i.e., $w_{j,s}\in\{0,1\}$. $\mathsf{AutoNN}$, detailed in Sec.˜2.3, uses weights to carefully pool together the benefits of $\mathsf{TSNN}$ and $\mathsf{DRNN}$. $\mathsf{AWNN}$ sadhukhan2025adaptivelyweightednearestneighborsmatrix improves upon $\mathsf{RowNN}$ by adaptively choosing the weights which optimally balances the bias-variance tradeoff of $\mathsf{RowNN}$ as follows
$\big(w_{1}^{\star}(i,t),...,w_{N}^{\star}(i,t)\big):=\mathop{\mathrm{argmin}}_{(v_{1},...,v_{N})\in\Delta_{N}}2\log(2N)\widehat{\sigma}^{2}\sum_{k\in[N]}v_{k}^{2}+\sum_{k\in[N]}v_{k}A_{k,t}\rho_{i,k}^{\mathrm{row}}.$ (6)
where $\widehat{\sigma}^{2}$ is the estimated error and $\Delta_{N}$ is a simplex in N; see sadhukhan2025adaptivelyweightednearestneighborsmatrix for details of ˜6. Tab.˜1 contains a concise summary of the existing nearest neighbor variants; see App.˜B for a detailed exposition for each methods.
@@
$n=1$ $\mathsf{RowNN}$ (li2019nearest) (Alg.˜1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{ColNN}$ (li2019nearest) (Alg.˜1) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq 0,\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{TSNN}$ (sadhukhan2024adaptivity) (Alg.˜2) $(x-y)^{2}$ $(x-y)^{2}$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq\eta_{2}})$
$\mathsf{AWNN}$ (sadhukhan2025adaptivelyweightednearestneighborsmatrix) (Alg.˜5) $(x-y)^{2}$ $(x-y)^{2}$ $w_{j}^{\star}(i,t)\cdot\mathbf{1}({\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{DRNN}$ (dwivedi2022doubly) (Alg.˜3) $\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$
$\mathsf{AutoNN}$ (Sec.˜2.3) $\alpha\cdot\mathsf{DRNN}+(1-\alpha)\cdot\mathsf{TSNN}$
$n>1$ $\mathsf{KernelNN}$ (choi2024learning) (Alg.˜4) $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(x,y)$ $\mathsf{MMD}_{\mathbf{k}}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
$\mathsf{W_{2}NN}$ (feitelberg2024distributional) (Alg.˜4) $\widehat{\mathsf{W}}_{2}^{2}(x,y)$ $\mathsf{W}_{2}^{2}(x,y)$ $\mathbf{1}({\rho_{i,j}^{\mathrm{row}}\leq\eta_{1},\rho_{s,t}^{\mathrm{col}}\leq 0})$
Under the distributional matrix completion setting ($n>1$ in ˜2), the methods $\mathsf{KernelNN}$ and $\mathsf{W_{2}NN}$ in Tab.˜1 take $\mu,\nu\in\mathcal{X}$ as square integrable probability measures, and $\varphi(\mu,\nu)$ as either the squared maximum mean discrepency (i.e. $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu,\nu)$, see muandet2017kernel) or squared Wasserstein metric (i.e., $\mathsf{W}_{2}(\mu,\nu)$, see bigot2020). Further, the entry-wise distance $\widehat{\varphi}(x,y)$ in this case is either the unbiased U-statistics estimator $\widehat{\mathsf{MMD}}_{\mathbf{k}}^{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{MMD}_{\mathbf{k}}^{2}(\mu_{i,t},\mu_{j,s})$ (see muandet2017kernel) or the quantile based estimator $\widehat{\mathsf{W}}_{2}(Z_{i,t},Z_{j,s})$ for $\mathsf{W}_{2}(\mu_{i,t},\mu_{j,s})$ (see bigot2020).